Let T: R3-R2 be the linear map given by a 2c (a) Find a basis of the range space. (Be sure to justify that it spans and is liPlease put the solution in the form of a formal proof, Thank You.
Answer #1


T:\mathbb{R}^3\rightarrow \mathbb{R}^2\\ T(a,b,c)=(a+2c,c)

First will find matrix of T,

T(1,0,0)=(1,0)=1(1,0)+0(0,1)--(1)

T(0,1,0)=(0,0)=0(1,0)+0(0,1)--(2)

T(0,0,1)=(2,1)=2(1,0)+1(0,1)--(3)

By(1)(2)(3)

[T]=\begin{bmatrix} 1 & 0&2 \\ 0 & 0& 1 \end{bmatrix}

Rref[T]=\begin{bmatrix} 1 & 0&0 \\ 0 & 0& 1 \end{bmatrix}

Rank of [T] is 2 .

And basis for Range(T) is \{(1,0),(0,1)\}

Null(T) is the set of solution space of [T]x=0

which is equivalent to

Rref[T]x=0\\ \begin{bmatrix} 1 & 0&0 \\ 0 & 0& 1 \end{bmatrix}\begin{bmatrix} x_1\\ x_2\\ x_3\end{bmatrix}=0\\ x_1=0,x_3=0\\ x=\begin{bmatrix} 0\\ 1\\ 0\end{bmatrix}

dim(Null(T))=1

Null(T)=Span\{\begin{bmatrix} 0\\ 1\\ 0\end{bmatrix}\}

Basis for Nll(T) is

010

(c)

Rank nullity theorem,

Rank(T) + Nullity(T) = Rank R 2+1=3